# Visual-Friendly Concept Protection via Selective Adversarial Perturbations

### v2

@@ V-E Hyper-parameter Analysis
Sensitivity Analysis of Training Iteration. Usually, the larger the training steps, the larger the value of $\mathcal{L}^{\prime}_{\theta}$ at different time steps, and the more successful the protection is. As shown in Fig. 14 (a), when increasing the number of training iterations, FID will first improve rapidly and then rise slowly, and the overall trend is that the longer the training iterations, the worse the quality of the protected image. Whereas LIQE will fall rapidly and then level off, the protection effect will improve and then keep stable. Finally, we selected 720 iterations based on the inflection point observed in Fig. 14 (a). At this point, we achieve substantial protection effectiveness (LIQE-SUM 4.37) while maintaining reasonable image quality (FID 27.04). Increasing iterations beyond 720 provides only marginal improvements in protection while continuing to degrade image quality.
Sensitivity Analysis of $c$. $c$ represents stronger constraints on perturbation visibility. As shown in Fig. 14 (b), increasing parameter $c$ reduces the FID score from 68.18 at $c=0.0$ to 15.33 at $c=1.0$, indicating significant improvements in visual quality of protected images. Meanwhile, protection effectiveness (LIQE-SUM) increases from  3.15 at $c=0.0$ to  5.43 at $c=0.5$, then slightly decreases to  4.75 at $c=1.0$. Based on the analysis, we selected $c=0.1$ as the optimal value, balancing acceptable protection effectiveness with maintainable visual quality.
@@
V-F Analysis of Adversary Settings
Following Anti-DreamBooth [8], we consider three settings in this paper: “convenient setting”, “adverse setting”, and “uncontrolled setting”.
@@ V-F Analysis of Adversary Settings
Convenient Setting. In a convenient setting (white-box setting), we (protectors) have all knowledge of the resources of adversary, including a pre-trained text-to-image generator, training term (e.g., “sks”), and training prompt. This scenario is practical because adversaries are likely to use high-quality, open-source pre-trained generators, with Stable Diffusion being the most prominent publicly available model. Additionally, people often rely on open-source code with default training terms and prompts. The convenient setting is the simplest for the protectors, and all relevant experiment results have been analyzed and mentioned in Sec. V-B.
TABLE IV: Results of model mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ denotes that higher values are preferable, while ↓ indicates the opposite. The best result in each column is highlighted in bold. SD is short for Stable Diffusion.
@@
TABLE V: Results of training term mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
TABLE VI: Results of training prompt mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
TABLE VII: Results of different post-processing with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
TABLE VIII: Results of different uncontrolled settings with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
  • Model Mismatching. In this scenario, protectors use Stable Diffusion v1-4 without knowing the specific version of Stable Diffusion employed by the adversary. Results in Fig. 8 and Table IV show effective facial area protection even if adversaries used Stable Diffusion v1-4/v1-5/v2-1, which reflects the transferability of VCPro.
  • Training Term Mismatching. We set protectors to use “sks” as a training term in Textual Inversion and DreamBooth, and adversaries use “t@t”. As shown in Fig. 8 and Table V, VCPro is robust to different training terms.
  • Training Prompt Mismatching. We set protectors to use “a photo of sks person.” as training prompt in Textual Inversion and DreamBooth, and adversaries to use “a dslr portrait of sks person.”. As shown in Fig. 8 and Table VI, VCPro can still successfully protect the important concepts.
  • Post-Processing. In this setting, we assume that the adversaries will use some post-processing to reduce the impact of the protected images, taking the two common post-processing techniques, Gaussian filtering and JPEG compression, as an example. Here we set the kernel size $7\times 7$ for Gaussian filtering and adopt 75% quality compression for JPEG, which also mentioned in [43] can be a good purification method. As shown in Fig. 8 and Table VII, although the protective effect is somewhat diminished, the final generated image is still visibly cracked. VCPro can still successfully protect the target area.
Uncontrolled Settings. In this section, we examine the scenario where an adversary has access to a set of original images of the target concept, which are then combined with protected images to train DreamBooth. We evaluate three settings in which the proportion of original images is varied from 25% to 75%, as detailed in Table VIII. All experiments employ a perturbation size of $\epsilon$ = 8/255 on the VGGFace2 dataset and utilize Stable Diffusion v1-4 as the base model. Our protection remains effective as long as the proportion of protected images exceeds 50%, but its effectiveness diminishes as the proportion of original images increases. Ideally, VCPro would receive platform support, become widely adopted, and be implemented across all social media platforms, thereby mitigating the risks associated with these uncontrolled settings.
@@
TABLE IX: Results of frequency analysis. The dataset is VGGFace2, and the perturbation size is $8/255$. ↑ denotes that higher values are preferable, while ↓ indicates the opposite. Average Low/High Frequency Difference is calculated based on the average L1 norm per pixel. The best result in each column is highlighted in bold.
@@
V-G Frequency Domain Analysis
We analyzed the frequency characteristics of VCPro via Discrete Wavelet Transform (DWT) with Haar wavelet decomposition, consistent with our perturbation generation process. This analysis is significant as human visual perception exhibits more sensitivity in low-frequency components than in high-frequency details. We compared two variants: VCPro with frequency constraint, and without frequency constraint where $D(\cdot)=||{x}-{x^{\prime}}||_{2}^{2}$ in Eq. 8. Further implementation details are provided in the supplementary materials.
@@ V-G Frequency Domain Analysis
Statistical Analysis in Different Frequency Components. Table IX provides a statistical analysis to quantify the frequency distribution of perturbations using the VGGFace2 dataset. With our frequency constraint, both low-frequency (6.80) and high-frequency (6.98) differences are substantially lower than without the constraint (17.39 and 11.47, respectively). Meanwhile, the frequency constraint effectively shifts perturbation distribution, reducing low-frequency components from 60.26% (unconstrained) to 49.37% (constrained). Both findings effectively explain the improved visual quality of the constrained approach as low-frequency perturbations are more perceptible to human vision.
VI Conclusion
Future Works and Limitations. We have not yet reached the capability to generate protected images in real time. However, we are committed to ongoing research and development in this area. Our future efforts will focus on finding efficient methods to produce protected images swiftly while maintaining high visual quality. By doing so, we aim to significantly enhance the overall user experience and ensure that our solutions meet the highest standards of both functionality and aesthetics.
Acknowledgments
We would like to express our gratitude to Yu Wu for his valuable discussions and support during the early stages of the experiments. Thanks are also due to Zijun Zhou for developing the user study system, and to Xiaoyu Kong for his assistance in the aesthetic aspects of the figures.
Appendix A Case Show
Appendix B Hyper-parameter Analysis
Sensitivity Analysis of Training Iteration. Usually, the larger the training steps, the larger the value of $\mathcal{L}^{\prime}_{\theta}$ at different time steps, and the more successful the protection is. As shown in Fig. 14 (a), when increasing the number of training iterations, FID will first improve rapidly and then rise slowly, and the overall trend is that the longer the training iterations, the worse the quality of the protected image. Whereas LIQE will fall rapidly and then level off, the protection effect will improve and then remain stable. Finally, we selected 720 iterations based on the inflection point observed in Fig. 14 (a). At this point, we achieve substantial protection effectiveness (LIQE-SUM 4.37) while maintaining reasonable image quality (FID 27.04), and Fig. 15 also shows similar pheonmentions. Increasing iterations beyond 720 provides only marginal improvements in protection while continuing to degrade image quality.
Sensitivity Analysis of $c$. $c$ represents stronger constraints on perturbation visibility. As shown in Fig. 14 (b), increasing parameter $c$ reduces the FID score from 68.18 at $c=0.0$ to 15.33 at $c=1.0$, indicating significant improvements in visual quality of protected images. Meanwhile, protection effectiveness (LIQE-SUM) increases from  3.15 at $c=0.0$ to  5.43 at $c=0.5$, then slightly decreases to  4.75 at $c=1.0$. Fig. 16 presents qualitative results across different $c$ values, demonstrating that smaller $c$ values (e.g., $c=0.0$) provide stronger protection but compromise visual quality, while larger $c$ values (e.g., $c=1.0$) lead to friendly visibility with reduced protection. Based on analysis, we selected $c=0.1$ as the optimal value, balancing acceptable protection effectiveness with maintainable visual quality.
@@ Appendix B Hyper-parameter Analysis
Appendix C Details of Frequency Analysis
@@
    Low Frequency $\text{Low Frequency Ratio}=\frac{\text{Mean}(\Delta L)}{\text{Mean}(\Delta L)+\text{Mean}(\Delta H)}\times 100\%$
    High Frequency $\text{High Frequency Ratio}=\frac{\text{Mean}(\Delta H)}{\text{Mean}(\Delta L)+\text{Mean}(\Delta H)}\times 100\%$

### v3

@@ Appendix D Hyper-parameter Analysis
Sensitivity Analysis of Training Iteration. Usually, the larger the training steps, the larger the value of $\mathcal{L}^{\prime}_{\theta}$ at different time steps, and the more successful the protection is. As shown in Fig. 6 (a), when increasing the number of training iterations, FID will first improve rapidly and then rise slowly, and the overall trend is that the longer the training iterations, the worse the quality of the protected image. Whereas LIQE will fall rapidly and then level off, the protection effect will improve and then keep stable. Finally, we selected 720 iterations based on the inflection point observed in Fig. 6 (a). At this point, we achieve substantial protection effectiveness (LIQE-SUM 4.37) while maintaining reasonable image quality (FID 27.04). Fig. 7 also shows similar pheonmentions. Increasing iterations beyond 720 provides only marginal improvements in protection while continuing to degrade image quality.
Sensitivity Analysis of $c$. $c$ represents stronger constraints on perturbation visibility. As shown in Fig. 6 (b), increasing parameter $c$ reduces the FID score from 68.18 at $c=0.0$ to 15.33 at $c=1.0$, indicating significant improvements in visual quality of protected images. Meanwhile, protection effectiveness (LIQE-SUM) increases from  3.15 at $c=0.0$ to  5.43 at $c=0.5$, then slightly decreases to  4.75 at $c=1.0$. Fig. 8 presents qualitative results across different $c$ values, demonstrating that smaller $c$ values (e.g., $c=0.0$) provide stronger protection but compromise visual quality, while larger $c$ values (e.g., $c=1.0$) lead to friendly visibility with reduced protection. Based on the analysis, we selected $c=0.1$ as the optimal value, balancing acceptable protection effectiveness with maintainable visual quality.
@@
Appendix E Analysis of Adversary Settings
Following Anti-DreamBooth (le_etal2023antidreambooth), we consider three settings in this paper: “convenient setting”, “adverse setting”, and “uncontrolled setting”.
@@ Appendix E Analysis of Adversary Settings
Convenient Setting. In a convenient setting (white-box setting), we (protectors) have all knowledge of the resources of adversary, including a pre-trained text-to-image generator, training term (e.g., “sks”), and training prompt. This scenario is practical because adversaries are likely to use high-quality, open-source pre-trained generators, with Stable Diffusion being the most prominent publicly available model. Additionally, people often rely on open-source code with default training terms and prompts. The convenient setting is the simplest for the protectors, and all relevant experiment results have been analyzed and mentioned in our Experiment Section.
Table 5: Results of model mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ denotes that higher values are preferable, while ↓ indicates the opposite. The best result in each column is highlighted in bold. SD is short for Stable Diffusion.
@@
Table 6: Results of training term mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
Table 7: Results of training prompt mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
Table 8: Results of different post-processing with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
Table 9: Results of different uncontrolled settings with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
@@
  • Model Mismatching. In this scenario, protectors use Stable Diffusion v1-4 without knowing the specific version of Stable Diffusion employed by the adversary. Results in Fig. 10 and Table 5 show effective facial area protection even if adversaries used Stable Diffusion v1-4/v1-5/v2-1, which reflects the transferability of VCPro.
  • Training Term Mismatching. We set protectors to use “sks” as a training term in Textual Inversion and DreamBooth, and adversaries use “t@t”. As shown in Fig. 10 and Table 6, VCPro is robust to different training terms.
  • Training Prompt Mismatching. We set protectors to use “a photo of sks person.” as training prompt in Textual Inversion and DreamBooth, and adversaries to use “a dslr portrait of sks person.”. As shown in Fig. 10 and Table 7, VCPro can still successfully protect the important concepts.
  • Post-Processing. In this setting, we assume that the adversaries will use some post-processing to reduce the impact of the protected images, taking the two common post-processing techniques, Gaussian filtering and JPEG compression, as an example. Here we set the kernel size $7\times 7$ for Gaussian filtering and adopt 75% quality compression for JPEG, which also mentioned in (sandoval2023jpeg) can be a good purification method. As shown in Fig. 10 and Table 8, although the protective effect is somewhat diminished, the final generated image is still visibly cracked. VCPro can still successfully protect the target area.
Uncontrolled Settings. In this section, we examine the scenario where an adversary has access to a set of original images of the target concept, which are then combined with protected images to train DreamBooth. We evaluate three settings in which the proportion of original images is varied from 25% to 75%, as detailed in Table 9. All experiments employ a perturbation size of $\epsilon$ = 8/255 on the VGGFace2 dataset and utilize Stable Diffusion v1-4 as the base model. Our protection remains effective as long as the proportion of protected images exceeds 50%, but its effectiveness diminishes as the proportion of original images increases. Ideally, VCPro would receive platform support, become widely adopted, and be implemented across all social media platforms, thereby mitigating the risks associated with these uncontrolled settings.
@@
Table 10: Results of frequency analysis. The dataset is VGGFace2, and the perturbation size is $8/255$. ↑ denotes that higher values are preferable, while ↓ indicates the opposite. Average Low/High Frequency Difference is calculated based on the average L1 norm per pixel. The best result in each column is highlighted in bold.
@@
Appendix F Frequency Domain Analysis
[…]
@@ Appendix F Frequency Domain Analysis
Statistical Analysis in Different Frequency Components. Table 10 provides a statistical analysis to quantify the frequency distribution of perturbations using the VGGFace2 dataset. With our frequency constraint, both low-frequency (6.80) and high-frequency (6.98) differences are substantially lower than without the constraint (17.39 and 11.47, respectively). Meanwhile, the frequency constraint effectively shifts perturbation distribution, reducing low-frequency components from 60.26% (unconstrained) to 49.37% (constrained). Both findings effectively explain the improved visual quality of the constrained approach as low-frequency perturbations are more perceptible to human vision.
@@
Appendix G Details of Frequency Analysis
@@
    $\text{Low Frequency Ratio}=\frac{\text{Mean}(\Delta L)}{\text{Mean}(\Delta L)+\text{Mean}(\Delta H)}\times 100\%$
    $\text{High Frequency Ratio}=\frac{\text{Mean}(\Delta H)}{\text{Mean}(\Delta L)+\text{Mean}(\Delta H)}\times 100\%$